# Instruct-Tuning Pretrained Causal Language Models for Ancient Greek Papyrology and Epigraphy

### v1

@@ Abstract
This article presents an experiment in fine-tuning a pretrained causal language model (Meta’s Llama 3.1 8B Instruct) for aiding in three fundamental tasks of philological research: chronological and geographic attribution as well as text restoration in ancient Greek inscriptions and documentary papyri. Using a prompt-based instruct approach, the fine-tuned models surpass the state of the art in key metrics. For inscriptions, the models achieve a lower average character error rate (CER) of 22.5% (vs. 26.3%), while closely matching top-1 accuracy (60.9% vs. 61.8%) and top-20 accuracy (77.5% vs. 78.3%) for sequences up to 10 characters. They also provide a practical advantage by ignoring spaces during reconstruction, aligning better with the scriptio continua format of ancient written artifacts. In geographic attribution, the model outperforms previous benchmarks with a top-1 accuracy of 75.0% (vs. 70.8%) and a top-3 accuracy of 83.7% (vs. 82.1%). For dating, it achieves an average deviation of 26.2 years (vs. 29.3) and a median deviation of 1 year (vs. 3) from the actual date range. The models also set new baselines for documentary papyri, with a CER of 16.3%, a top-1 accuracy of 71.3%, and top-20 of 85.0% in text reconstruction; a top-1 accuracy of 66.4% and top-3 of 79.9% in geographic attribution; and, in chronological attribution, a deviation of 21.7 years from the actual termini post/ante quem, with a median deviation of 0 years.
Keywords papyrology  $\cdot$ epigraphy  $\cdot$ philology  $\cdot$ deep learning  $\cdot$ large language models  $\cdot$ instruct models
@@ 1 Introduction
A significant number of documentary texts in ancient Greek—such as announcements, laws, receipts, contracts, and letters—have survived from the Archaic period through to the early Byzantine era. These texts were inscribed or written on various materials, often on durable stone blocks or sheets of papyrus, which have survived particularly well in drier climates.
@@
A key task of philological research is to assign dates to these documents and attribute them to their places of origin. Insights can typically be drawn from the provenance of the artifact in question, its physical features and writing style, but typically the analysis also relies on clues from the textual content, such as the information conveyed or specific linguistic features. Another important task is the reconstruction of damaged or missing letters in the fragments, which are frequently marred by gaps, abrasions, and scribal errors. Grammatical and stylistic criteria, as well as parallels from similar documents—now readily searchable through databases like www.papyri.info and inscriptions.packhum.org—can be invoked to support or challenge specific conjectures or emendations. However, the conception of such suggestions has traditionally relied on the linguistic creativity and inventiveness of individual scholars, and on their personal expertise acquired over many years of reading and working with similar ancient texts. Conjectural “higher" criticism is performed ope ingenii (“with the aid of one’s ingenuity”), as textual critics put it. It is a branch of philological research that often seems to align more closely with the realm of art than with that of science, or at least with a scientific field where, to borrow a phrase from T.S. Eliot [1, p. 103], “there is no method except to be very intelligent.”
@@
This traditional division within textual criticism—between the objective, technical analysis of manuscript data and the subjective, intuitive insight of an astute critic—is undergoing a significant transformation. Machine learning has the potential to contribute across all areas of philology, including those traditionally characterized by interpretation, guesswork, and creativity. Language models can be pretrained on the entire surviving textual record of ancient civilizations and fine-tuned to parse, classify, and analyze various aspects of words, phrases, and texts in these languages. They can be applied to explore relationships between manuscripts, identify prosodic, rhetorical, and narrative patterns, detect intertextual relationships, recognize themes, analyze sentiments, summarize ideas, and translate texts (see review by Sommerschield and colleagues [2]). More immediately relevant to our purposes, they can be trained to date and geographically attribute texts, detect errors, generate corrections, and fill in missing words and phrases ([3, 4, 5, 6, 7]).
@@
In the field of Greek documentary texts, Assal and colleagues [4] remain the leading pioneers. They trained a transformer model with a sparse attention mechanism (BigBird), combining word- and character-level tokenization, to date, attribute geographically, and reconstruct damaged sequences in ancient Greek inscriptions. On their test set of 7,811 inscriptions, they reported the following results:
@@
  • Text Restoration: The model achieved a top-1 accuracy of 61.8% for restoring sequences of 1 to 10 characters, compared to 25.3% for the human experts they recruited. The character error rate (CER) for the model was 26.3%, while the CER for the philologists was 59.6%. Additionally, 78.3% of the correct cases were found within the top 20 results.
@@
  • Chronological Attribution: The model’s predictions were, on average, 29.3 years away from the actual dating span (termini post and ante quem), with a median error of only 3 years.
The modest aim of this experiment was to advance research in machine learning and textual criticism in three respects:
First, the model published by Assael and colleagues [4] counts spaces as characters in the restoration task. For example, the sequence ϰαι ο λογ – – – – – – ος τον ϑεον would not be decoded as ϰαι ο λογος ην προς τον ϑεον. Instead, the input would need to be ϰαι ο λογ – – – – – – – – ος τον ϑεον, forcing us to guess beforehand that damaged characters 3 and 6 are empty spaces. In ancient inscriptions and papyri—typically written in scriptio continua or at least without evenly marked spaces—we often know the approximate number of missing letters but not the combined number of missing characters and word boundaries. This study aims to develop a model that tracks only characters, better reflecting the real-world situations philologists face when working with damaged textual artifacts.
Second, given the rapid advancements in deep learning and Natural Language Processing over the last two years, it is worth exploring how newer, larger pretrained causal language models—such as Meta’s LLaMA, Google’s Gemma, Microsoft’s Phi, or OpenAI’s GPT-4o—can assist textual critics. One promising approach could involve enabling masked language modeling in these models and retraining them to fill in missing sequences of text within a bidirectional context (see the first step of the recipe developed by [8], LLM2Vec). However, in this preliminary study, I have opted for a much simpler approach that is immediately applicable and open to experimentation with larger and newer models in the future: using a straightforward instruct template prompting the model to generate the date, location, or reconstruction at the end of a sequence.
Third, I aimed to expand the application of deep learning for textual criticism of ancient Greek documentary texts to include papyri in addition to inscriptions. While Pavlopoulos and colleagues [9] developed a regression model that dates Greek papyri with an average error of 54 years, there has otherwise been little strictly text-based deep-learning research on documentary papyri.
@@ 2 Method
Date spans, geographic attributions, and texts for the ancient Greek inscriptions were sourced from the Packard Humanities Institute database (inscriptions.packhum.org/) using the text processing pipeline developed and shared by Assael and colleagues [4] (github.com/sommerschield/iphi). The texts of Greek documentary papyri were obtained from the Duke Databank of Documentary Papyri (DDbDP), with their chronological and geographical metadata sourced from the Heidelberger Gesamtverzeichnis der griechischen Papyrusurkunden Ägyptens (HGV), both accessible via https://papyri.info.
@@
To facilitate comparison with the results reported by Assael and colleagues [4], the datasets were cleaned of diacritics. Commas were removed and all other punctuation marks converted to a high dot “·”. The data was then split into a train (95%) and test sets (5%).
@@
Dates for both papyri and inscriptions were formatted as exact integers corresponding to the midpoint of their respective date spans. For dates with only a terminus post quem, they were set 25 years after the given date, while those with only a terminus ante quem were set 25 years before. Place name attributions for the papyri were normalized semi-manually due to inconsistencies across database entries. Although more detailed processing would be beneficial in the future, the current approach – admittedly rough (see https://github.com/ericu9500/PapyriAndInscriptions/train_data/data/places_and_dates.tsv) – was deemed sufficient for the experimental proof-of-concept purposes of this study.
All text editions were included in two versions: one in which philologists’ reconstructions of lost or uncertain characters were integrated, and another where these characters were replaced with “-”. Individual missing characters were represented by a single “-”, while longer stretches of missing text were indicated by 10 consecutive “-”.
@@
Text masks were generated before tokenization by randomly selecting spans of 3 to 20 characters, ensuring that no more than 50% of a sequence of intact characters was masked. Longer spans of intact characters were favored by assigning weights proportional to the square of their lengths, increasing the probability of selecting longer spans for masking. These spans were replaced with a simple placeholder in the format “[6 letters missing]”, where the count excluded spaces and punctuation marks within the masked text segment.
@@
The training data was formatted using the same chat template that the instruct model had been pretuned with, utilizing one of three system prompts: “Date this papyrus fragment/inscription to an exact year!”, “Assign this papyrus fragment/inscription to an exact place!”, or “Reconstruct the missing letters in this papyrus fragment/inscription!”. The user input consisted of a version of the text, while the assistant’s response provided the corresponding date, place, or masked sequence of characters, including spaces and punctuation marks. The text was tokenized using the model’s TikToken-based tokenizer.
@@
The model, LLaMA 3.1 8B-Instruct [10], was fine-tuned on 4 A100 GPUs with 40 GB and batch size of 8 or 80 GB and batch size of 20 depending on availability at the time, utilizing the toolkit Torchtune (github.com/pytorch/torchtune). In an initial calibration round, exploratory fine-tuning was conducted with three different data combinations to identify the best approach: one model trained solely on papyri (4 epochs, approximately 80 hours), another solely on inscriptions (4 epochs, approximately 120 hours), and a third on both inscriptions and papyri (3 epochs, approximately 132 hours). For each data combination, two versions were trained: one using the entire sequence of system, user, and assistant messages, and another with the prompt masked out for loss computation.
@@
Subsequently, the best-performing models were further trained separately for the three tasks until their performance plateaued on the test set. During this stage, entries shorter than 75 tokens or longer than 847 tokens were filtered out, and data augmentation was applied by randomly altering the order of sentences and adding noise through the replacement of 5%, 10%, 15%, 20%, and 25% of preserved characters with “-”. Also, masked sequences of 1–20 (and not only 3-20) characters were included. In the final stage, the fine-tuned models were re-merged with the base model using the TIES method developed by Yadav and colleagues [11] with the toolkit Mergekit (github.com/arcee-ai/mergekit).
@@
The models were tested using the same metrics proposed by Assael and colleagues [4]: CER (Character Error Rate), top-1 and top-20 accuracy for normalized reconstructions (spaces, punctuation marks and numerals ignored), top-1 and top-3 accuracy for geographical attributions, and average and median deviation from the date ranges provided in the database. Entries shorter than 90 characters were excluded. For inscriptions, the test was conducted on 7,811 samples for restoration, 2,870 samples for geographical attribution, and 1,856 samples for dating. For papyri, the test was conducted on 7,811 samples for restoration, 1,990 samples for geographical attribution, and 2,295 samples for dating. Dates, places, and reconstructions were evaluated based on top-1, top-3, and top-20 rankings, respectively, out of 45 beams. For reconstructions, a repetition penalty of 1.2 was applied to encourage varied predictions.
@@ 3.1 Calibration
For the dating task, again, training with the prompt masked out led to better performance, with an average difference of 16 years for inscriptions and 3 years for papyri after 4 epochs. The transfer effect of training the same model on both papyri and inscriptions was slightly positive for papyri (-2 years) but negative for inscriptions (0.3 years), after 3 epochs.
@@
  • Reconstruction: Three epochs for inscriptions (CER 28.1%, top-1 accuracy 52.8%, and top-20 accuracy 68.4%) and four epochs for papyri (CER 21.0%, top-1 accuracy 64.9%, and top-20 accuracy 78.3%).
@@
  • Dating: Three epochs for inscriptions (average distance 40.5 years, median distance 5 years) and three epochs (trained on both papyri and inscriptions) for papyri (average distance 22.11 years, median distance 0 years).
@@ 3.1.1 Continued training
For the dating task, performance plateaued after 1 additional epoch (for a total of 4 epochs) for the inscriptions, with an average deviation of 26.2 years and a median deviation of 1 year from the actual dating span. Training on the papyrus dataset for dating also plateaued after 1 additional epoch (for a total of 4 epochs), with an average deviation of 21.7 years and a median deviation of 0 years from the actual dating span.
@@ 3.1.2 Re-merging
In the final step, the effects of re-merging the fine-tuned models with the original base model were assessed. The TrIm, Elect Sign & Merge method (TIES-Merging), where parameters that changed only slightly during fine-tuning are trimmed ([11]), was applied, using the toolkit Mergekit (github.com/arcee-ai/mergekit). This process led to a deterioration of 10% in geographical attribution accuracy for inscriptions and 5% for papyri, along with an increase in the average error for dating by 4 years for inscriptions and 3 years for papyri.

### v2

# Instruction-Tuning Pretrained Causal Language Models for Text Restoration, Chronological Attribution, and Geographic Attribution of Ancient Greek Papyri and Inscriptions

@@ Abstract
This article presents an experiment in fine-tuning a pretrained causal language model (Meta’s Llama 3.1 8B Instruct) to assist with three key tasks in philological research: dating, localizing, and restoring missing or illegible characters in ancient Greek inscriptions and documentary papyri.
Utilizing a straightforward instruction-based approach and a 95%/5% train/test split, the papyri models achieved a character error rate (CER) of 14.9%, a top-1 accuracy of 73.5%, and a top-20 accuracy of 86.0% in text reconstruction. For geographic attribution, they reached a top-1 accuracy of 66.4% and a top-3 accuracy of 79.9%. In chronological attribution, the models demonstrated an average deviation of 21.7 years from the actual terminus post/ante quem, with a median deviation of 0 years.
For inscriptions, the models achieved a CER of 20.5%, a top-1 accuracy of 63.7%, and a top-20 accuracy of 83.0% for sequences up to 10 characters. In geographic attribution, they attained a top-1 accuracy of 75.0% and a top-3 accuracy of 83.7%, while in dating, they had an average deviation of 37.1 years and a median deviation of 3 years from the actual date range.
Benchmarked against the state-of-the-art model (Ithaca) on a shared test set and on recently edited inscriptions, the instruction-tuned models excelled in text restoration, while also offering the practical advantage of ignoring spaces during reconstruction, which aligns with the scriptio continua format of ancient texts. However, their performance in geographic and chronological attribution was lower than Ithaca’s.
The results are preliminary but suggest that fine-tuning larger pretrained causal language models using instruction templates for philological research holds significant promise, particularly in the area of textual criticism.
Keywords papyrology  $\cdot$ epigraphy  $\cdot$ philology  $\cdot$ deep learning  $\cdot$ large language models  $\cdot$ instruction models
@@ 1 Introduction
A significant number of documentary texts in ancient Greek – such as announcements, laws, receipts, contracts, and letters – have survived from the Archaic period through to the early Byzantine era. These documents were inscribed or written on various materials, often on durable blocks of stone or sheets of papyrus, which have been preserved primarily in drier climates.
@@
A key task of philological research is to assign dates to these texts and attribute them to their places of origin. Insights can typically be drawn from the provenance of the artifact in question, its physical features, and writing style. However, the analysis also heavily relies on clues from the textual content itself, such as the information conveyed or specific linguistic features. Another important challenge is the reconstruction of damaged or missing letters in these fragments, which are frequently marred by gaps, abrasions, and scribal errors. Grammatical and stylistic criteria, as well as parallels from similar documents—now readily searchable through databases such as www.papyri.info and inscriptions.packhum.org—are often invoked to support or challenge specific conjectures or emendations. Traditionally, such suggestions have been the product of the linguistic creativity and inventiveness of individual scholars, drawing on expertise accumulated over years of engaging with similar ancient texts. Conjectural “higher” criticism is performed ope ingenii (“with the aid of one’s ingenuity”), as textual critics put it. It is a branch of philological research that seems to align as much with the realm of art as with science—or, at least, with a scientific field where, to borrow a phrase from T.S. Eliot [1, p. 103], “there is no method except to be very intelligent.”
@@
This traditional division within textual criticism—between the objective, technical analysis of manuscript data and the subjective, intuitive insight of an astute critic—is undergoing a significant transformation. Machine learning has the potential to contribute across all areas of philology, including those traditionally characterized by interpretation, conjecture, and creativity. Language models can be pretrained on the entire surviving textual record of ancient civilizations and fine-tuned to parse, classify, and analyze various aspects of words, phrases, and texts in these languages. They can be applied to explore relationships between manuscripts, identify prosodic, rhetorical, and narrative patterns, detect intertextual connections, recognize themes, analyze sentiments, summarize ideas, and translate texts (see review in [2]). More directly relevant to our purposes, they can be trained to date and geographically attribute texts, detect errors, generate corrections, and fill in missing words and phrases ([3, 4, 5, 6, 7]).
@@
In the field of Greek documentary texts, Assael and colleagues [4] were and remain the leading pioneers. They trained a transformer model with a sparse attention mechanism (BigBird), combining word- and character-level tokenization, to date, attribute geographically, and reconstruct damaged sequences in ancient Greek inscriptions. On their test set of 7,811 inscriptions, they reported the following results:
@@
  • Text Restoration: The model achieved a top-1 accuracy of 61.8% for restoring sequences of 1 to 10 characters, compared to 25.3% for the human experts they recruited. The character error rate (CER) for the model, averaged over character lengths 1–10, was 26.3%, while the corresponding figure for philologists was 59.6%. Additionally, in 78.3% of the cases, the correct reconstruction was found within the top 20 results.
@@
  • Chronological Attribution: The model’s calculated dates, derived from the probability distribution over each century between 800 BCE and 800 CE, were on average 29.3 years off from the actual (termini post and ante quem), with a median error of only 3 years.
The aim of this experiment is to advance research in machine learning and textual criticism in three specific respects:
First, the model published by Assael and colleagues [4] counts spaces as characters in the restoration task. For example, the sequence ϰαι ο λογ – – – – – – ος τον ϑεον would not be decoded as ϰαι ο λογος ην προς τον ϑεον. Instead, the input would need to be ϰαι ο λογ – – – – – – – – ος τον ϑεον, requiring the researcher to guess beforehand that damaged characters 3 and 6 are empty spaces. In ancient inscriptions and papyri—typically written in scriptio continua or at least without clearly marked spaces—we often know the approximate number of missing letters but not the precise number of missing characters and word boundaries. This study aims to develop a model that tracks only characters, better reflecting the real-world situations philologists face when working with damaged textual artifacts.
Second, given the rapid advancements in deep learning and natural language processing over the last two years, it is worth exploring how newer, larger pretrained causal language models—such as Meta’s LLaMA, Google’s Gemma, Microsoft’s Phi, or OpenAI’s GPT-4o—can assist textual critics. One promising approach could involve enabling masked language modeling in these models and retraining them to fill in missing sequences of text within a bidirectional context (see the first step of the recipe developed by [8], LLM2Vec). However, in this preliminary study, I have opted for a simpler approach that is immediately applicable and open to experimentation with larger and newer models in the future: using a straightforward instruct template, prompting the model to generate the date, location, or reconstruction at the end of the sequence after a special token.
Third, I aim to expand the application of deep learning for textual criticism of ancient Greek documentary texts to include papyri in addition to inscriptions. While Pavlopoulos and colleagues [9] developed a regression model that dates Greek papyri with an average error of 54 years, there has otherwise been little strictly text-based deep-learning research on documentary papyri.
@@ 2 Method
Date spans, geographic attributions, and texts for the ancient Greek inscriptions were sourced from the Packard Humanities Institute database (inscriptions.packhum.org/) using the text processing pipeline developed and shared by Sommershield and colleagues [10] (github.com/sommerschield/iphi). The texts of Greek documentary papyri were obtained from the Duke Databank of Documentary Papyri (DDbDP), with their chronological and geographical metadata sourced from the Heidelberger Gesamtverzeichnis der griechischen Papyrusurkunden Ägyptens (HGV), both accessible via https://papyri.info’s GitHub repository at https://github.com/papyri.
@@
The datasets were cleaned of diacritics, commas were removed, and all other punctuation marks were converted to a high dot, “·”. The data was then split into training (95%) and test sets (5%).
@@
Dates for both papyri and inscriptions were formatted as discrete numbers corresponding to the midpoint of their respective date spans. For dates with only a terminus post quem, they were set 25 years after the given date, while those with only a terminus ante quem were set 25 years earlier. Place name attributions for the papyri were normalized semi-manually due to inconsistencies across database entries. While more detailed processing would be beneficial in the future, the current approach – though admittedly rough (see https://github.com/ericu9500/PapyriAndInscriptions/train_data/data/places_and_dates.tsv) – was deemed sufficient for the proof-of-concept purposes of this study. All text editions were included in two versions: one in which philologists’ reconstructions of lost or uncertain characters were integrated, and another where these characters were replaced with “-”. Individual missing characters were represented by a single “-”, while longer stretches of missing text were indicated by 10 consecutive hyphens.
@@
Text masks were generated prior to tokenization by randomly selecting spans of 3 to 20 characters, ensuring that no more than 50% of a sequence of intact characters was masked. Longer spans of intact characters were favored by assigning weights proportional to the square of their lengths, increasing the probability of selecting longer spans for masking. These spans were replaced with a simple placeholder in the format “[6 letters missing]”, where the count excluded spaces and punctuation marks within the masked text segment.
@@
The training data was formatted using the same chat template that the instruction model had been pre-tuned with, utilizing one of three system prompts: “Date this papyrus fragment/inscription to an exact year!”, “Assign this papyrus fragment/inscription to an exact place!”, or “Reconstruct the missing letters in this papyrus fragment/inscription!”. The user input consisted of a version of the text, while the assistant’s response provided the corresponding date, place, or masked sequence of characters, including spaces and punctuation marks. The text was tokenized using the model’s TikToken-based tokenizer.
@@
The base model, LLaMA 3.1 8B-Instruct [11], was fine-tuned on 4 A100 GPUs with 40 GB of memory and a batch size of 8, or on 80 GB GPUs with a batch size of 20, depending on availability at the time, using the toolkit Torchtune (github.com/pytorch/torchtune). In an initial calibration round, exploratory fine-tuning was conducted with three different data combinations to identify the best approach: one model trained solely on papyri (4 epochs, approximately 80 hours), another solely on inscriptions (4 epochs, approximately 120 hours), and a third on both inscriptions and papyri (3 epochs, approximately 132 hours). For each data combination, two versions were trained: one using the entire sequence of system, user, and assistant messages, and another with the prompt masked out for loss computation.
@@
Subsequently, the best-performing models were further trained separately for the three tasks until their performance plateaued on the test set. During this stage, entries shorter than 75 tokens or longer than 847 tokens were filtered out, and data augmentation was applied by randomly altering the order of sentences and adding noise through the replacement of 5%, 10%, 15%, 20%, and 25% of preserved characters with “-”. Additionally, masked sequences of 1–20 characters (not only 3–20) were included. In the final stage, the fine-tuned models were re-merged with the base model using the TIES method developed by Yadav and colleagues [12], utilizing the Mergekit toolkit (github.com/arcee-ai/mergekit).
@@
The models were evaluated using the same metrics proposed by Assael and colleagues [4]: CER (Character Error Rate) averaged across lengths 1–10, top-1 and top-20 accuracy for normalized reconstructions (where spaces, punctuation marks, and numerals were ignored), top-1 and top-3 accuracy for geographical attributions, and average and median deviation from the date ranges provided in the database. Entries shorter than 90 characters were excluded. For inscriptions, the test was conducted on 7,811 samples for restoration, 2,870 samples for geographical attribution, and 1,856 samples for dating. For papyri, the test was conducted on 7,811 samples for restoration, 1,990 samples for geographical attribution, and 2,295 samples for dating. Dates, places, and reconstructions were evaluated based on top-1, top-3, and top-20 rankings, respectively, out of 60 beams.
In an additional step, the models were fine-tuned for text restoration on texts truncated to 50–750 characters to allow for comparison with the model developed by Assael and colleagues [4]. They were tested on a subset of my test set that coincided with theirs (entries with a PHI ID ending in 3). This subset included 264 samples with dates, 435 samples with places, and 4,110 masked samples based on 411 papyri for restoration (100 samples of each character length). Additionally, the models were evaluated on 2 previously unedited Uppsala papyri (1,620 masked samples) and 6 recently edited inscriptions (4,111 masked samples) to assess their performance on texts that could not have been included in the pretraining data of Meta’s base model. Different subpages of both http://papyri.info and https://inscriptions.packhum.org sporadically appear in the indices of the Common Crawl corpus (http://index.commoncrawl.org).
@@ 3.1 Calibration
For the dating task, again, training with the prompt masked out led to better performance, with an average difference of 16 years for inscriptions and 3 years for papyri after 4 epochs. The transfer effect of training the same model on both papyri and inscriptions was slightly positive for papyri (-2 years) but negative for inscriptions (0.3 years) after 3 epochs.
@@
  • Reconstruction: Three epochs for inscriptions (CER 27.9%, top-1 accuracy 52.2%, and top-20 accuracy 71.0%) and four epochs for papyri (CER 20.6%, top-1 accuracy 64.0%, and top-20 accuracy of 80.86%).
@@
  • Dating: Three epochs for inscriptions (average distance 40.3 years, median distance 5 years) and three epochs (trained on both papyri and inscriptions) for papyri (average distance 22.11 years, median distance 0 years).
@@ 3.1.1 Continued training
For the dating task, performance plateaued after 1 additional epoch (for a total of 4 epochs) for the inscriptions, with an average deviation of 37.1 years and a median deviation of 3 years from the actual dating span. Training on the papyrus dataset for dating also plateaued after 1 additional epoch (for a total of 4 epochs), with an average deviation of 21.6 years and a median deviation of 0 years from the actual dating span.
@@ 3.1.2 Re-merging
In the final step, the effects of re-merging the fine-tuned models with the original base model were assessed. The TrIm, Elect Sign & Merge method (TIES-Merging), where parameters that changed only slightly during fine-tuning are trimmed ([12]), was applied, using the toolkit Mergekit (github.com/arcee-ai/mergekit). This process led to a deterioration of 10% in geographical attribution accuracy for inscriptions and 5% for papyri, along with an increase in the average error for dating by 4 years for inscriptions and 3 years for papyri. For the reconstruction tasks, however, there was only insignificant deterioration of 0.1–0.2% in CER and top-1 accuracy, but a 3% increase in top-20 accuracy. Hence, the model’s overall restoration capacity interestingly improved when it was re-infused with the base model.
3.1.3 (Imperfect) comparison with Ithaca
An accurate comparison with the state-of-the-art model for these tasks developed by Assael and colleagues [4] – Ithaca – is not possible, since the present model was trained for experimental purposes with a 95%/5% split as opposed to Ithaca’s 80/10/10. Yet some form of comparison is still useful. Accordingly, in a final step, the models were finetuned on the dataset where every entry was truncated to 50–750 characters (minimum and maximum length of entries for Ithaca). The resulting inscription restoration model and Ithaca were then run on two test sets: one solely derived from entries within my 5% test set that coincided with Assael and colleagues’ test split (where PHI ID ends in 3), and one derived from inscriptions that were not in the PHI or openly available on the internet as of December 2023 (4111 samples based on 6 inscriptions; see appendix). On the shared test set, the finetuned Llama model had an average CER of 22.55% (vs. Ithaca’s 27.0%), a top-1 accuracy of 61.4% (vs. 60.9%), and top-20 accuracy of 77.00% (vs. 71.6%). On the new inscriptions, it had a CER of 30.8% (vs. Ithaca’s 31.1%), a top-1 accuracy of 48.9% (vs. 46.5%), and a top-20 accuracy of 72.9% (vs. 68.3%).
In geographic attribution of inscriptions in the shared test set, the Llama model had a top-1 accuracy of 65.0% (vs. Ithaca’s 69.0%) and top-20 accuracy of 78.4% (vs. Ithaca’s 80.2%). In chronological attribution, it had an average distance of 55.7 (vs. Ithaca’s 48.3) and a median distance of 4 (vs. Ithaca’s 4.5).
The Papyrus model on the recent papyri showed variability: in the case of Pap.Ups 106, it had a CER of 21.0%, top-1 accuracy of 62.5%, and top-20 accuracy of 79%. With Pap.Ups. 18, it had a CER of 38.6%, top-1 accuracy of 37.2%, and top-20 accuracy of 61.2%. It plausibly assigns the former document to Oxyrhynchus and accurately dates it to 71 AD (the fourth year of Vespasian, 71–72 AD); equally plausibly, it places the latter in Oxyrhynchus and dates it to 304 AD, approximately 11–12 years from the correct date of around 292 AD (the eighth year of Diocletian and the seventh of Maximian are mentioned).
These further-tuned models were also tested on the entire 5% test set with good results. The inscription model had a final CER of 20.5%, a top-1 accuracy of 63.7%, and top-20 accuracy of 83.0%. The Papyrus model had a CER of 14.9%, top-1 accuracy of 73.5%, and top-20 accuracy of 85.9%. Accordingly they were chosen as the final models for restoration.